OpenStudy (timaashorty):

I'm preparing for the ACT , and forgot how to do this. Can anyone help me? What is the slope of a line passing through r(2,-6) and S(-4,-3) in the (x,y) coordinate plane?

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ R&({\color{red}{ 2}}\quad ,&{\color{blue}{ -6}})\quad S &({\color{red}{ -4}}\quad ,&{\color{blue}{ -3}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}}\)

OpenStudy (timaashorty):

\[\frac{(-3) -1 }{ (-4)- 2 } = \frac{ -4 }{ -6} = \frac{ 2 }{ 3 } ?\] But then 2/3 is not in my answer choices

OpenStudy (timaashorty):

ooh is my mistake i put -1 instead of -6 for y^1

OpenStudy (timaashorty):

\[\frac{ -3 - (-6)}{ -4 - 2 } = \frac{ 3 }{ -6 } = - \frac{ 1 }{ 2 }\] *

OpenStudy (timaashorty):

Thanks for refreshing my mind ! (:

OpenStudy (jdoe0001):

yw